# Blackbox Dataset Inference for LLM

### v2

@@ 5.1. Overview
We propose a method leveraging reference models to realize the idea inspired by Observation II. The method involves three steps to determine if suspect model $\mathcal{M}$ used victim dataset $\mathcal{D}$ for training.
  1. (1)
@@
    We collect a set of non-member reference models that are never trained on $\mathcal{D}$, annotated as $\mathbf{R}$ = $\{\mathcal{R}_{1},\mathcal{R}_{2},...,\mathcal{R}_{n}\}$. We fine-tune each reference model with $\mathcal{D}$ to produce member reference models $\mathbf{R}^{t}$ = $\{\mathcal{R}_{1}^{t},\mathcal{R}_{2}^{t},...,\mathcal{R}_{n}^{t}\}$.
  2. (2)
@@
    We pre-process $\mathcal{D}$ to identify candidates of tainted samples, based on the similarity between the oracle outputs and the responses from the original and trained reference models.
  3. (3)
@@
    We compare $\mathcal{M}$ with original reference models and trained reference models on tainted samples, determining if $\mathcal{M}$ used $\mathcal{D}$ according to which reference set $\mathcal{M}$ is closer to.
@@
For easy reference, important annotations used in the rest of this paper are summarized in Table 3.
Table 3. Important notations used in this paper.
Notations Descriptions
$\mathcal{M}$ The suspect model
$\mathcal{D}$ The victim dataset
$\mathbf{R}$ The reference models never trained on $\mathcal{D}$
$\mathbf{R}^{t}$ The reference models fine-tuned on $\mathcal{D}$
$\mu$ The size threshold used to pre-filter samples
$\mathbf{O}$ The responses from $\mathbf{R}$ to samples passing pre-filtering
$\mathbf{O}^{t}$ The responses from $\mathbf{R}^{t}$ to samples passing pre-filtering
$\delta^{t}$ The similarity disparity threshold in identifying tainted samples
$\delta^{s}$ The similarity disparity threshold in dataset inference
5.2. Building Reference Models
Initial Building: Building the reference models is straightforward. Given victim dataset $\mathcal{D}$, we identify a set of $n$ open-source models following several criteria. First, we ensure the models did not use $\mathcal{D}$ in training. The easiest way is to focus on models released before $\mathcal{D}$. If that is infeasible, we opt for models that attach the list of datasets they are trained on (e.g.,  (jiang2023mistral, ; team2024gemma, ; llama3modelcard, ; qwen2, ; glm2024chatglm, )) and pick those not using $\mathcal{D}$. Second, we configure and run the models in their instruct mode to make sure they can respond to inputs from $\mathcal{D}$. We do not consider their base models because the model is usually released to the market in its instruct mode. Third, we recommend models with a middle size (e.g., 5B - 10B parameters). This way, we have models that are knowledgeable enough but remain resource-friendly. By default, we use five reference models, including Mistral-7B-v0.1 (jiang2023mistral, ), gemma-7b (team2024gemma, ), Meta-Llama-3-8B (llama3modelcard, ), Qwen2-7B (qwen2, ), glm-4-9b (glm2024chatglm, ). All the models are representative and widely used.
@@ 5.2. Building Reference Models
Fine-tuning: To establish the trained version of reference models, we rely on fine-tuning since we are out of access to their pre-training process. Considering that full-parameter fine-tuning (FPFT) (rajbhandari2020zero, ; rasley2020deepspeed, ) is both time-consuming and resource-heavy, we adopt parameter efficient fine-tuning (PEFT) (lora, ). By default, we use QLORA (dettmers2024qlora, ), a representative PEFT method with quantization. The hyper-parameters we used to fine-tune the models are discussed in §6.1.
5.3. Identifying Tainted Samples
Our method is built on top of tainted samples, which in general cause the suspect model to dramatically change its responses before and after being involved in training. We leverage the reference models, both trained and non-trained, to identify candidates of tainted samples.
@@ 5.3. Identifying Tainted Samples
Pre-filtering: Responses with a small size do not offer sufficient information for making a meaningful decision. Thus, we filter out samples that receive a response shorter than $\mu$ bytes from $\mathcal{M}$. We set $\mu$ to 20 by default. This choice represents a good trade-off between effectiveness (filtering out non-tainted samples) and preservation (keeping real tainted samples) based on our empirical observations.
@@
Selection: Each sample passing pre-filtering is then processed for identifying tainted samples. We feed its input to the original reference models, obtaining their responses $\mathbf{O}$ = $\{\mathcal{O}_{1},\mathcal{O}_{2},...,\mathcal{O}_{n}\}$. Similarly, we get the responses from the fine-tuned reference models $\mathbf{O}^{t}$ = $\{\mathcal{O}_{1}^{t},\mathcal{O}_{2}^{t},...,\mathcal{O}_{n}^{t}\}$. For each pair of $\mathcal{O}_{i}$ and $\mathcal{O}_{i}^{t}$ ($1\leq i\leq n$), we respectively compute their BERTScore with the oracle output, deriving a pair of similarity score ($\mathcal{S}_{i}$, $\mathcal{S}_{i}^{t}$). We consider the sample tainted if:
(1) $\forall i\in[1,n],\,\mathcal{S}_{i}^{t}-\mathcal{S}_{i}>\delta^{t}$
@@
Briefly, the above condition requires that the responses from every reference model, before and after training on $\mathcal{D}$, present a disparity of $\delta^{t}$ in similarity score. We perceive it as a good strategy because it indicates the training incurs prevalent variations in model responses. The hyper-parameter $\delta^{t}$ controls the balance between recall (how many true tainted samples are identified) and precision (how many false tainted samples are included). We set it to 30% by default to prefer recall. Based on our observations, a higher threshold can result in fewer samples left with certain datasets, making dataset inference infeasible.
5.4. Dataset Inference
With the candidate tainted samples, we infer whether $\mathcal{M}$ used $\mathcal{D}$ for training. As clarified in Observation II, we aim to determine if the amount of tainted samples identifiable with $\mathcal{M}$ is closer to the situation where $\mathcal{D}$ is involved in training or to the other scenario where $\mathcal{D}$ is absent. Yet, we only have access to $\mathcal{M}$ in one situation, not both. To this end, we rely on the two sets of reference models, alternatively inspecting which set $\mathcal{M}$ is closer to.
@@ 5.4. Dataset Inference
Processing Tainted Samples: For each tainted sample we identified above, we inspect whether $\mathcal{M}$ responds closer to the non-trained reference models or their trained versions. The process is similar to selecting tainted samples. We get the responses from each pair of reference model and its trained version, annotated as $\mathcal{O}_{i}$ and $\mathcal{O}_{i}^{t}$. Respectively calculating their BERTScore similarity with the response of $\mathcal{M}$, we obtain their similarity scores with $\mathcal{M}$, represented as $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{t}$. We consider $\mathcal{M}$ responds closer to the trained reference set if:
(2) $\forall i\in[1,n],\,\mathcal{S}_{i}^{t}-\mathcal{S}_{i}>\delta^{s}$
@@
where $n$ represents the number of reference models. For simplicity, we call the sample a positive tainted sample. As noted, we require $\mathcal{M}$ to act closer to every trained reference model to determine a positive tainted sample. This is intended to reduce randomness. The pre-configured parameter $\delta^{s}$ controls the confidence of decisions, and it defaults to the value of $\delta^{t}$. If the above condition does not hold, we consider $\mathcal{M}$ responds closer to the non-trained reference set, and we call the sample a negative tainted sample.
@@
To mitigate the potential randomness in a single response, we obtain multiple responses from the suspect model for the same question and calculate a $\delta^{s}$ for each response. We use the largest $\delta^{s}$ to make decisions. By default, we obtains three responses for each question.
@@
Inferring Model Membership: If we observe more positive tainted samples than negative ones from $\mathcal{D}$, we consider $\mathcal{M}$ used $\mathcal{D}$ for training (i.e., a member model). Otherwise, we consider it did not (a non-member model).
6. Evaluation
6.1. Experimental Setup
@@
Table 4. Datesets used in our study. “Part” explains which split from the official dataset is used in our experiments.
@@
databricks-dolly-15k (Databricks_dataset, ) databricks train 15,011
alpaca (Alpaca_dataset, ) tatsu-lab train 52,002
SlimOrca (Slimorca_dataset, ) Open-Orca train 517,982
OpenHermes-2.5 (Openhermes_dataset, ) teknium train 1,001,551
@@
Target Datasets: To evaluate the performance of our method, we collect victim datasets from Hugging Face following several criteria. ➀ The datasets are frequently downloaded and used by the public, offering representativeness. ➁ The datasets are formatted as input-output pairs such that our method can be applied without further subjective data processing. ➂ The datasets have been used in training a group of publicly released models. This enables us to collect member models in the wild and, thus, assess our method in realistic settings. After broad searching, we identify four target datasets, including databricks-dolly-15k, alpaca, SlimOrca, and OpenHermes-2.5. As summarized in Table 4, they have different owners and sizes. Due to the intensive amount of experiments and limited GPU budget, we randomly sample 50,000 data points for our evaluation if the dataset is larger than that.
@@
Suspect Models: We prioritize suspect models from the wild to better mirror the reality. Datasets released on Hugging Face comes with a card showing the list of public models trained or fine-tuned on them (e.g., (Databricks_dataset, )). Based on that, we identify 10 member models that can properly follow instructions for each target dataset, and we call them wild member models. Similarly, we collect models released on Hugging Face but not included in a dataset’s card to work as non-member models. Non-member models are more prevalent and we identify 20 of them for each target dataset. Besides checking their capability of following instructions, we also ensure they can understand English corpora. Details of the member and non-member models are summarized in Table 10.
Table 5. Local models used in our evaluations.
Owner Name
mistralai Mistral-7B-v0.3 (mistral7Bv3, )
google gemma-2b (gemma2B, )
meta-llama Llama-3.1-8B (llama3.1_8B, )
meta-llama Llama-3.2-3B (llama3.2_3B, )
Qwen Qwen1.5-4B (qwen, )
Qwen Qwen1.5-7B (qwen, )
Qwen Qwen2.5-3B (qwen2.5, )
Qwen Qwen2.5-7B (qwen2.5, )
THUDM glm-2b (DBLP:conf/acl/DuQLDQY022, )
THUDM glm-10b (DBLP:conf/acl/DuQLDQY022, )
@@
To increase the number of member models to balance with the non-member models, we further prepare 10 local member models for each dataset atop widely-used base models presented in Table 5. Specifically, we fine-tune the base models with QLORA (qlora, ) on each target dataset to create new member models. The parameters of QLORA we adjusted include rank of matrices (8), scaling factor (32), dropout probability (0.05), bias type (none), optimizer (paged_adamw_8bit), learning rate ($1\times 10^{-4}$), batch size (8), training epoch (3) [^1] Smaller models need more training epochs to reach acceptable training loss. Thus, for gemma-2b and glm-2b, we set their training epoch to 4., target modules (all non-linear hidden layers). All other parameters are set to their default values.
@@
Reference Models: We build our reference models on top of five popular base models, including Mistral-7B-v0.1 (jiang2023mistral, ), gemma-7b (team2024gemma, ), Meta-Llama-3-8B (llama3modelcard, ), Qwen2-7B (qwen2, ), and glm-4-9b (glm2024chatglm, ). Specifically, we take their corresponding instruct models (Mistral-7B-Instruct-v0.1, gemma-7b-it, Meta-Llama-3-8B-Instruct, Qwen2-7B-Instruct, and glm-4-9b-chat)—which can better follow instructions—as the non-member reference models $\mathbf{R}$. We further fine-tune the base models on the target dataset to produce the member reference models $\mathbf{R}^{t}$. The fine-tuning follows the same procedure as training our local suspect models.
@@
Dataset Inference (Non-IID): We apply our method, leveraging the reference models above, to the 20 wild non-member models, 10 wild member models, and 10 local member models on each dataset. We measure the recall, precision, and F1 score of the results [^2] Recall = $\frac{TP}{TP+FN}$, Precision = $\frac{TP}{TP+FP}$, F1-Score = $2\times\frac{Precision*Recall}{Precision+Recall}$, where TP, FP, FN represent true positives, false positives, and false negatives..
@@
Dataset Inference (IID): In the evaluation above, the non-member models are trained on data irrelevant to the target dataset. Maini et al. (maini2024llm, ) pointed out a more challenging situation where the training of the non-member models involves a dataset that is independent and identically distributed (IID) with respect to the target dataset.
We extend an evaluation to simulate the IID situation. Specifically, we evenly split a target dataset $\mathcal{D}$ into $\mathcal{D}_{x}$ and $\mathcal{D}_{y}$ and de-duplicate highly similar samples between them [^3] We exclude a data sample from $\mathcal{D}_{y}$ if its input and output both have a 0.8+ Jaccard similarity (jaccard, ) with another sample in $\mathcal{D}_{x}$.. $\mathcal{D}_{x}$ is considered as the victim dataset and adopted to prepare the reference models. The wild non-members presented in Table 10 are fine-tuned on $\mathcal{D}_{y}$ to work as non-member models. We further fine-tune the five reference models on $\mathcal{D}_{y}$ to obtain extra non-member models. This aims to assess the cases where the reference models and the non-members overlap and, thus, carry higher detection difficulties. In this evaluation, we only focus on detecting non-members, as the IID setting makes no difference to members.
@@
Baseline: To the best of our knowledge, there are no existing methods for black-box dataset inference. Yet, DPDLLM (dpdllm, ) offers black-box membership inference of individual data samples against LLMs, which we adapt to provide dataset inference.
@@
In practical blackbox settings, the suspect model is private, preventing the acquisition of member samples and non-member samples. We bypass this restriction in our evaluation as follows. If the suspect model is a member model (i.e., trained on the target dataset), we randomly pick 1,000 samples from the target dataset as member samples. Otherwise, we fine-tune the suspect model with another dataset and pick 1,000 random samples from that dataset. To obtain non-member samples, we randomly select 1,000 samples from a dataset not used in training the suspect model.
@@
Table 6. Dataset inference results under the non-IID setting. “HF” refers to Hugging Face. A value “x/y” indicates that “x” out of “y” members/non-members are correctly identified.
@@
6.2. Dataset Inference Accuracy
Non-IID: The dataset inference results under Non-IID settings are summarized in Table 6. Our method presents 100% accuracy, regardless of the target dataset and the suspect model. Fundamentally, training with $\mathcal{D}$ leads the target model to behaving closer to the trained reference models on tainted samples. As shown in the top half of Figure 4, this enables our method to see a larger group of positive tainted samples and identify the model as a member. In contrast, the target model—without training on $\mathcal{D}$—diverges from the trained reference models on tainted samples. Thus, our method reports most tainted sample to be negative and clearly detects the model as a non-member, as shown in the bottom half of Figure 4.

### v3

@@ V-A Overview
We propose a method leveraging reference models to realize the idea inspired by Observation II. The method involves three steps to determine if the suspect model $\mathcal{M}$ used the victim dataset $\mathcal{D}$ for training.
  1. 1.
@@
    We collect a set of non-member reference models that are never trained on $\mathcal{D}$, annotated as $\mathbf{R}$ = $\{\mathcal{R}_{1},\mathcal{R}_{2},...,\mathcal{R}_{n}\}$. We then fine-tune each reference model with $\mathcal{D}$ to produce member reference models $\mathbf{R}^{t}$ = $\{\mathcal{R}_{1}^{t},\mathcal{R}_{2}^{t},...,\mathcal{R}_{n}^{t}\}$.
  2. 2.
@@
    We pre-process $\mathcal{D}$ to identify candidates of tainted samples, based on the similarity between the oracle outputs and the responses from the non-member and member reference models.
  3. 3.
@@
    We compare the outputs of $\mathcal{M}$ with non-member and member reference models on tainted samples, determining if $\mathcal{M}$ used $\mathcal{D}$ according to which reference set $\mathcal{M}$ is closer to.
@@
For easy reference, important annotations used in the rest of this paper are summarized in Table VIII, and pseudo-code of our method is presented in Algorithm 1.
1
2
Input: Victim Dataset $\mathcal{D}$; Suspect Model $\mathcal{M}$; Non-Member Reference Models $\mathbf{R}=\{\mathcal{R}_{1},...,\mathcal{R}_{n}\}$; Thresholds $\mu$, $\delta^{t}$, $\delta^{s}$
Output: Whether $\mathcal{M}$ used $\mathcal{D}$ during fine-tuning
3
// (1) Build Member Reference Models
4
5for $i=1$ to $n$ do
6    Fine-tune $\mathcal{R}_{i}$ on $\mathcal{D}$ via QLoRA to obtain $\mathcal{R}_{i}^{t}$;
7
8 end for
[…]
10
// (2) Identify Tainted Samples
11
[…]
13
14foreach $(x,y)\in\mathcal{D}$ do
15    if $|y|<\mu$ then
16       continue
17    end if
18
19   $\Delta$ = get_diff($\mathbf{R}$, $\mathbf{R}^{t}$, $x$, $y$, $n$)
   // All diff must be larger than the threshold
20    if count($\Delta>\delta^{t}$) == $n$ then
21        $T\leftarrow T\cup\{(x,y)\}$;
22
23    end if
24
25 end foreach
26
// (3) Compare $\mathcal{M}$ with Reference Models
27
[…]
29
30foreach $(x,y)\in T$ do
31
   // $\mathcal{M}(x)$ is the response from $\mathcal{M}$ given $x$
32    $\Delta_{min}^{i}\leftarrow$ min(get_diff($\mathbf{R}$, $\mathbf{R}^{t}$, $x$, $\mathcal{M}(x)$, $n$)) ;
    // Sampling $k$ times to reduces randomness
33    $\Delta_{\max}$ = max($\{\Delta_{min}^{1},...,\Delta_{min}^{k}\}$);
34
35   if $\Delta_{\max}>\delta^{s}$ then
36        $pos\leftarrow pos+1$;
37
38   else
39        $neg\leftarrow neg+1$;
40
41    end if
42
43 end foreach
44
// (4) Final Decision
45
46if $pos>neg$ then
47    return $\mathcal{M}$ used $\mathcal{D}$;
48
49else
50    return $\mathcal{M}$ did not use $\mathcal{D}$;
51
52 end if
53
54
551ex Function get_diff($\mathbf{R}$, $\mathbf{R}^{t}$, $x$, $y$, $n$):
56    $\Delta=\emptyset$;
57    for $i=1$ to $n$ do
58        $\mathcal{S}_{i}\leftarrow\text{BERTScore}(\mathcal{R}_{i}(x),y)$;
59        $\mathcal{S}_{i}^{t}\leftarrow\text{BERTScore}(\mathcal{R}_{i}^{t}(x),y)$;
60
61       $\Delta\leftarrow\Delta\cup\{\mathcal{S}_{i}^{t}-\mathcal{S}_{i}\}$
62    end for
63   return $\Delta$
64
Algorithm 1 Black-Box Dataset Inference with Tainted Samples
V-B Building Reference Models
Initial Building: Building the reference models is straightforward. Given victim dataset $\mathcal{D}$, we identify a set of $n$ open-source models as non-member reference models. We ensure the models did not use $\mathcal{D}$ in training. The easiest way is to focus on models released before $\mathcal{D}$. If that is infeasible, we opt for models that attach the list of datasets they were trained on and pick those not using $\mathcal{D}$. By default, we use five reference models.
@@ V-B Building Reference Models
Fine-tuning: To establish the member reference models, we rely on fine-tuning since we are out of access to their pre-training process (lines 1–4 in Algorithm 1). Considering that full-parameter fine-tuning (FPFT) [32, 33] is both time-consuming and resource-heavy, we adopt parameter efficient fine-tuning (PEFT) [34]. By default, we use QLORA [35], a PEFT method with quantization. The hyper-parameters we used to fine-tune the reference models are discussed in §VI-A.
V-C Identifying Tainted Samples
@@
Pre-filtering: Responses with a short length do not offer sufficient information for making a meaningful decision (e.g. Yes/No, Multiple-choices). Thus, we filter out samples whose oracle outputs are shorter than $\mu$ tokens (lines 7–9). We set $\mu$ to 20 by default. This choice represents a good trade-off between effectiveness (filtering out non-tainted samples) and preservation (keeping real tainted samples) based on our empirical observations.
@@
Selection: The samples passed pre-filtering are then processed for identifying tainted samples (lines 10–13). We feed its input to the non-member reference models, obtaining their responses $\mathbf{R}(x)$ = $\{\mathcal{R}_{1}(x),\mathcal{R}_{2}(x),...,\mathcal{R}_{n}(x)\}$. Similarly, we get the responses from the member reference models $\mathbf{R}^{t}(x)$ = $\{\mathcal{R}^{t}_{1}(x),\mathcal{R}^{t}_{2}(x),...,\mathcal{R}^{t}_{n}(x)\}$. For each pair of $\mathcal{R}_{i}$ and $\mathcal{R}_{i}^{t}$ ($1\leq i\leq n$), we respectively compute their BERTScore with the oracle output, deriving a pair of similarity score ($\mathcal{S}_{i}$, $\mathcal{S}_{i}^{t}$). We consider the sample is tainted if:
$\forall i\in[1,n],\,\mathcal{S}_{i}^{t}-\mathcal{S}_{i}>\delta^{t}$ (1)
@@
We perceive it as a good strategy because it indicates the fine-tuning incurs prevalent variations in model responses. The hyper-parameter $\delta^{t}$ controls the balance between recall (how many true tainted samples are identified) and precision (how many false tainted samples are included). We set it to 0.3 by default. Based on our empirical observations shown in Table IX, using a higher threshold leaves too few samples for reliable analysis, whereas using a lower threshold introduces non-strict tainted samples. Both cases render dataset inference infeasible.
V-D Dataset Inference
@@
Processing Tainted Samples: For each tainted sample we identified above, we inspect whether $\mathcal{M}$ responds closer to the non-member reference models or their respective members (lines 15–24). The process is similar to selecting tainted samples. We get the responses from each pair of non-member and member reference models, annotated as $\mathcal{R}_{i}(x)$ and $\mathcal{R}^{t}_{i}(x)$. Respectively calculating their BERTScore similarity with the response of $\mathcal{M}$, we obtain their similarity scores with $\mathcal{M}$, represented as $\mathcal{S}_{i}$ and $\mathcal{S}_{i}^{t}$. We consider $\mathcal{M}$ responds closer to the member reference set if:
$\forall i\in[1,n],\,\mathcal{S}_{i}^{t}-\mathcal{S}_{i}>\delta^{s}$ (2)
@@
where $n$ represents the number of reference models. For simplicity, we call the sample a positive tainted sample if Equation 2 is satisfied. As noted, we require $\mathcal{M}$ to act closer to every member reference model to determine a positive tainted sample. Otherwise, we consider $\mathcal{M}$ responds closer to non-member reference set, and we call the sample a negative tainted sample. By default, we set $\delta^{s}$ as same as $\delta^{t}$.
@@
To mitigate the potential randomness in a single response, we obtain multiple responses from the suspect model for the same question and calculate a similarity score for each response (line 18). We use the largest similarity score to make decisions. By default, we obtains three responses for each question.
@@
Inferring Model Membership: If we observe more positive tainted samples than negative ones from $\mathcal{D}$, we consider $\mathcal{M}$ used $\mathcal{D}$ in fine-tuning (i.e. member model). Otherwise, we consider it did not (i.e. non-member model) (lines 25–29).
VI Evaluation
VI-A Experimental Setup
@@
TABLE III: Datesets used in our study. “Part” explains which split from the official dataset is used in our experiments.
@@
databricks-dolly-15k [26] databricks train 15,011
alpaca [36] tatsu-lab train 52,002
SlimOrca [37] Open-Orca train 517,982
OpenHermes-2.5 [38] teknium train 1,001,551
@@
Target Datasets: To evaluate the performance of our method, we collect four victim datasets from Hugging Face, including databricks-dolly-15k, alpaca, SlimOrca, and OpenHermes-2.5. As summarized in Table III, they have different owners and sizes. Due to the intensive amount of experiments and limited GPU budget, we randomly sample 50,000 data points for our evaluation if the dataset is larger than that.
@@
Suspect Models: We prioritize suspect models from the wild to better mirror the reality. Datasets released on Hugging Face comes with a card showing the list of public models fine-tuned on them (e.g., [26]). Based on that, we identify 10 member models called Hugging Face member models. Similarly, we collect models released on Hugging Face but not included in a dataset’s card to work as non-member models. Non-member models are more prevalent and we identify 20 of them for each target dataset. Details of the member and non-member models are summarized in LABEL:tab:app:models.
@@
To increase the number of member models to balance with the non-member models, we further prepare 10 local member models for each dataset atop widely-used models presented in Table X. Specifically, we fine-tune the models with QLORA [35] on each target dataset to create new member models. The parameters of QLORA we adjusted include rank of matrices (8), scaling factor (32), dropout probability (0.05), bias type (none), optimizer (paged_adamw_8bit), learning rate ($1\times 10^{-4}$), batch size (8), training epoch (3) [^1] Smaller models need more training epochs to reach acceptable training loss. Thus, for gemma-2b and glm-2b, we set their training epoch to 4., target modules (all non-linear hidden layers). All other parameters are set to their default values. Due to the limited computational resources, we use these hyperparameters instead of those suggested in [35]. However, these differences do not affect the results of our method. We obtain the same detection results using the suggested parameters by [35] as shown in Figure 10.
@@
Reference Models: We build our reference models on top of five popular models, including Mistral-7B-v0.1 [39], gemma-7b [40], Meta-Llama-3-8B [41], Qwen2-7B [42], and glm-4-9b [43]. Specifically, we take their corresponding instruct models which can better follow instructions as the non-member reference models $\mathbf{R}$. We further fine-tune the non-member models on the target dataset to produce the member reference models $\mathbf{R}^{t}$. The fine-tuning follows the same procedure as training our local member models.
@@
Dataset Inference (Non-IID): We apply our method, leveraging the above reference models, to detect non-member and member models on each dataset. We measure the recall, precision, and F1 score of the results. [^2] Recall = $\frac{TP}{TP+FN}$, Precision = $\frac{TP}{TP+FP}$, F1-Score = $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where TP, FP, and FN represent true positives, false positives, and false negatives, respectively.
@@
Dataset Inference (IID): In the evaluation above, the non-member suspect models are trained on data irrelevant to the victim dataset. Maini et al. [9] pointed out a more challenging situation where the training of the non-member models involves a dataset that is independent and identically distributed (IID) with respect to the victim dataset.
We extend our evaluation to simulate an IID setting. Specifically, we evenly split the original dataset $\mathcal{D}$ into two subsets, $\mathcal{D}{x}$ and $\mathcal{D}{y}$, and remove highly similar cross-subset samples [^3] A sample in $\mathcal{D}{y}$ is excluded if both its input and output exhibit a Jaccard similarity greater than 0.8 [44] with any sample in $\mathcal{D}{x}$.. We treat $\mathcal{D}{x}$ as the victim dataset, which is used to train both the member reference models and the member suspect models. The subset $\mathcal{D}{y}$, drawn from the same distribution as $\mathcal{D}_{x}$, is used to train non-member suspect models.
In this evaluation, detecting members can be omitted because the member-detection workflow and target are identical in both IID and non-IID settings. For non-member detection, we fine-tune the Hugging Face models listed in LABEL:tab:app:models on $\mathcal{D}{y}$ to construct non-member suspect models. Additionally, we fine-tune the five reference models on $\mathcal{D}{y}$ to obtain extra non-member suspect models. This setup allows us to examine scenarios where the reference models and suspect models highly overlap, thereby increasing the difficulty of distinguishing them.
@@
Baseline: To the best of our knowledge, there are no existing methods for black-box dataset inference. Yet, DPDLLM [12] and MIAVLM [13] offers black-box membership inference of individual data samples against LLMs and VLMs, which we adapt to provide dataset inference.
@@
MIAVLM leverages the variance in temperature sensitivity between member and non-member data for detection. Specifically, it probes the target model at varying temperatures and measure the rate of response divergence. Member sets exhibit significantly higher sensitivity, manifesting as rapid similarity degradation compared to non-members. These dynamic, set-level spectral features are utilized to train a inference classifier.
In practical blackbox settings, the suspect model is private, preventing the acquisition of member samples and non-member samples. We bypass this restriction in our evaluation as follows. If the suspect model is a member model (i.e., trained on the victim dataset), we randomly pick 1,000 samples from the victim dataset as member samples. Otherwise, we fine-tune the suspect model with another dataset and pick 1,000 random samples from that dataset. To obtain non-member samples, we randomly select 1,000 samples from a dataset not used in training the suspect model.
@@
TABLE IV: Dataset inference results under the non-IID setting. HF refers to models from Hugging Face. A value x/y indicates that x models are correctly identified as members/non-members among y models.
@@
VI-B Dataset Inference Accuracy
Non-IID: The dataset inference results under Non-IID settings are summarized in Table IV. Our method presents 100% accuracy, regardless of the victim dataset and the suspect model. Fundamentally, training with $\mathcal{D}$ leads the suspect model to behaving closer to the member reference models on tainted samples. As shown in the top half of Figure 3, this enables our method to see a larger group of positive tainted samples and identify the model as a member. In contrast, the suspect model—without training on $\mathcal{D}$—diverges from the member reference models on tainted samples. Thus, our method reports most tainted sample to be negative and clearly detects the model as a non-member, as shown in the bottom half of Figure 3.